OpenStudy (this):

-2x^2+8x-8=0 a, 0;two real solution b, 128; two real solution c,0; 2 imaginary solution d,0;one real solution

Parth (parthkohli):

Hint: you could start by dividing both sides by -2.

jhonyy9 (jhonyy9):

divide both sides by -2 and will get , ... ??? can you make it sure ...!

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (this):

yes

jhonyy9 (jhonyy9):

so than can you divide it ?

OpenStudy (this):

yes

jhonyy9 (jhonyy9):

and will get ---- ?

OpenStudy (this):

i don't if it is right or not but i got -8x-4=0

OpenStudy (anonymous):

on simplifying you get -x^2+4x-4=0 -x^2+2x+2x_4=0 now factorize: (-x+2)(-x+2)=0

jhonyy9 (jhonyy9):

so if you divide both sides by (-2) so than will result -2x^2 +8x -8 =0 /(-2) x^2 -4x +4=0 right ?

jhonyy9 (jhonyy9):

so what factorized will be eaxctly (x-2)^2 (x-2) squared right ?

jhonyy9 (jhonyy9):

so than x-2=0 x=2 than will be right d. so one real solution hope so much that is understandably sure right i wait your message bye

OpenStudy (anonymous):

yes

jhonyy9 (jhonyy9):

yes @akkib89 there you have missed anything ok ?

OpenStudy (anonymous):

actually it will be mod(x-2)=0

OpenStudy (anonymous):

-2x^2+8x-8=0 x^2 -4x +4 = 0 b^2-4ac 16 - 16 = 0 So..They have two equal Roots.....So..it means that....One SolN